OpenStudy (anonymous):

Which of the following describes the best method of solving two-step equations? A. Apply the multiplication property of equality first, and then apply the addition property of equality to isolate the variable. B. Apply the addition property of equality first, and then apply the multiplication property of equality to isolate the variable. C. Apply only the multiplication property of equality to isolate the variable. D. Apply only the addition property of equality to isolate the variable

OpenStudy (yanasidlinskiy):

\(\Huge\bf \color{yellow}{ Welcome~to~OpenStudy!!}\hspace{-310pt}\color{cyan}{Welcome~to~OpenStudy!!}\hspace{-307.1pt}\color{midnightblue}{Welcome~to~\color{purple}{Open}}\color{blue}{Study!!!!}\) Have you solved one of these before?

OpenStudy (anonymous):

no

OpenStudy (yanasidlinskiy):

Here's an example: \(\huge{-3x-7=-12}\) We need to get rid of the 7 since it's not with the variable. So instead subtracting, we add. And now we get, \(\huge{-3x-5}\) Divide both sides by -3 and you get... \(\huge{x=5/3}\) or, \[1\frac{ 2 }{ 3 }\]

OpenStudy (yanasidlinskiy):

Sorry, it took a little while.